# Eigenstate Thermalization Hypothesis (ETH) for off-diagonal matrix elements in integrable spin chains

### v1

@@ III.1.1 One-spin operator
In Fig. 3 we report $M_{ij}$ (cf. (17)) for the operator $E^{(11)}_{i}$ (cf. (16)). In Fig. 3 (a) we show the histograms $P(|M_{ij}|)$ for several chain sizes up to $L=240$, at fixed particle density $M/L=1/4$. The data are the same as in Fig. 2 (b) for the infinite-temperature macrostate. We observe that as $L$ increases, the center of the peak moves towards larger values of $|M_{ij}|$, and the distribution broadens. Moreover, the histograms are not symmetric around the maximum, suggesting non Gaussian behavior. Notice that at $L=240$ the typical off-diagonal matrix elements is $M_{ij}\sim 10^{-200}$. Indeed, to correctly capture such tiny matrix elements we employed arbitrary-precision arithmetic in evaluating the ABA formulas (see Appendix A). Let us now discuss the statistics of the off-diagonal matrix elements, i.e., the functional form of the distribution $P(M_{ij})$. For the Lieb-Liniger gas, Ref. [26] showed that $P(|{M}_{ij}|/L)$ is well described by a Fréchet distribution $P_{\alpha,\beta,\nu}$, whose probability distribution function takes the form
@@
$P_{\mu,\beta}(x)=\frac{1}{\beta}\,e^{-\frac{x-\mu}{\beta}-e^{-\frac{x-\mu}{\beta}}},$ (19)
The continuous lines in Fig. 3 are fits to (19) with $\beta,\mu$ the fitting parameters. For all the values of $L$ the data are quite well described by (19). To proceed, we observe that according to ETH (cf. (1)), one should expect that $M_{ij}\sim-L$, which has been verified in Ref. [26] for the Lieb-Liniger model. Thus, in the inset of Fig. 3 we report the histograms $P(M_{ij}/L)$. The collapse of the data for different $L$ is not perfect, as a residual “drift” towards the right is visible. This is expected, and it was observed also for the Lieb-Liniger gas [26]. This effect can be fully accounted for by a residual $L$ dependence of the parameter $\mu$ of the distribution, which is related to the average of the Gumbel distribution $\mu_{G}=\mu+\beta\gamma$, with $\gamma$ the Euler-Mascheroni constant. Thus, let us consider the average $\overline{M_{ij}}$ as a function of $L$. Following Ref. [26] we fit $\overline{M_{ij}}/L$ to the behavior
@@
with $c,c_{0}$ fitting parameters. By fitting the data with $L\geq 80$ we obtain $c\approx 0.13$ and $c_{0}\approx 0.82$. Now, the histograms of $|M_{ij}|-c\ln(L)$ for different values of $L$ should collapse on the same curve, at least in the limit $L\to\infty$. In Fig. 3 (b) we show the distribution of $|M_{ij}|/L-c\ln(L)$, with $c$ the fitted value. The data collapse for different $L$ improves as compared with Fig. 3 (a), although it is not perfect. This could be attributed to residual finite $L$ corrections. In the inset of Fig. 3 (b) we report $|\overline{M_{ij}}|/L$ versus $L$ (notice the logarithmic scale on the $x$-axis). The dashed line is the fit to (20). The crosses denote the average of the Gumbel distribution $\mu_{G}$ computed from the fitted constants $\mu,\beta$ extracted in panel (a), and are compatible with $\overline{M}_{ij}/L$. It is also interesting to consider the Cumulative Distribution Function (CDF). The CDF $F(x)$ of the Gumbel distribution is
[…]
@@
Eq. (21) implies that the double logarithm of $F(x)$ is a straight line. In Fig. 3 (c) we show the double logarithm of the CDF of $\left|M_{ij}\right|/L-c\ln(L)$ as extracted from the numerical data. The data exhibits an approximate linear behavior as a function $|M_{ij}|/L-c\ln(L)$ in the central region, whereas deviations from linear scaling are visible at the edges, which correspond to the tails of the distribution, where finite $L$ effects are expected to be larger.
@@
Figure 5: Distribution of $M_{ij}$ in Eq. 17 of the two-spin operator $E_{i}^{(11)}E_{i+1}^{(22)}$ for $L=20,24,32,40,$ and $48$ and $M=L/4$. (a) Histogram of the absolute value $|M_{ij}|$. The continuous lines are fits to the Gumbel distribution (19), with $\mu,\beta$ fitting parameters. (b) Histograms of $\left|M_{ij}\right|/L-c\ln(L)$, with $c$ obtained by fitting the average $\overline{M_{ij}}/L$ to the expected scaling (20). The fit is reported in the inset with the gray-dashed line. We report with the cross symbols the average of the Gumbel distributions $\mu_{G}=\mu+\beta\gamma$, with $\mu,\beta$ obtained from the fits in panel (a), and $\gamma$ the Euler-Mascheroni constant. (c) Double logarithm of the CDF of $\left|M_{ij}\right|/L-c\ln(L)$.
@@
An important question is how universal is the statistics of off-diagonal matrix elements. To investigate this aspect, we consider the infinite-temperature thermal ensemble at fixed particle density $M/L=1/2$. Numerical data are reported in Fig. 4. The qualitative behavior is similar to that at $M/L=1/4$ (see Fig. 3). The continuous lines in Fig. 4 (a) are fits to the Gumbel distribution (19), and are in good agreement with the numerical data. In the inset of Fig. 4 (a) we report the rescaled histograms of $\left|M_{ij}\right|/L$. The collapse is not perfect, similar to the case with $M/L=1/4$. We employ the same strategy as for $M/L=1/4$, i.e., fitting the average $\overline{M_{ij}}/L$ to (20), and plotting in Fig. 4 the shifted $M_{ij}/L-c\ln(L)$. The fit to the expected scaling (20) gives $c\approx 0.16,c_{0}\approx 0.94$ (see the inset in panel (b)). It is clear that the probability distribution $P({M}_{ij}-c\ln(L))$ is different than that for $M/L=1/4$. The data collapse for different $L$ is better than for $M/L=1/4$, and deviations near the peak become smaller upon increasing $L$. A fit of the histogram for $L=80$ to the Gumbel (19) gives $\beta\approx 0.29,\mu\approx 0.77$. Finally, in Fig. 4 (c) we report the double logarithm of the numerical CDF for $\left|M_{ij}\right|/L-c\ln L$. As for $M=1/4$, in the central region the CDF exhibits a clear linear behavior, as expected.
@@ III.1.2 Two-spin operator
Let us now discuss the statistics of off-diagonal matrix elements of local operators that have support on two nearest-neighbor sites. In the following we consider the infinite-temperature macrostate at $M/L=1/4$. Moreover, since we employ the approach of Ref. [25], as stressed already we are restricted to the macrostate constructed from eigenstates of the $XXX$ chain that do not contain bound states. Moreover, since the computational cost for evaluating the ABA formulas (see appendix A) increases exponentially with the size of the operator support, we provide results for $L\leq 60$. We anticipate that, despite this limitations, the qualitative scenario outlined for the one-spin operators remains the same. Let us focus on the operator $E_{i}^{(11)}E_{i+1}^{(22)}=(\mathds{1}_{i}-\sigma^{z}_{i})(\mathds{1}_{i+1}+\sigma^{z}_{i+1})/4$, which is sensitive to anti-aligned spin configurations on nearest-neighbor sites.
In Fig. 5 (a) we report the histogram of $M_{ij}$. The continuous lines are fits to the Gumbel distributions (19), which are in perfect agreement with the numerical data. In Fig. 5 (b) we show the histograms of the shifted $M_{ij}-c\ln(L)$, with $c$ obtained by fitting the average $\overline{M_{ij}}/L$ to (20) (see the inset in panel (b)). The fit gives $c\approx 0.25,c_{0}\approx 0.15$. Despite the smaller chain sizes as compared with Figs. 3 and 4, the data exhibit a reasonable collapse. A fit to the Gumbel distribution (19) gives $\beta\approx 0.28,\mu\approx 0$. Finally, in Fig. 5 (c) we plot the double logarithm of the CDF of $\widetilde{M}_{ij}$. Again, the approximate linear behavior in the central region confirms that $M_{ij}$ are well described by the Gumbel distribution.
@@ III.1.3 Non-Gaussianity of off-diagonal matrix elements
Our results for off-diagonal matrix elements confirm that, despite integrability, the finite-size scaling is similar to the ETH scaling (1), even quantitatively. Precisely, the exponential decay as $e^{-L}$ is the same, except for the correction (20). The difference between integrable and non-integrable models emerges strikingly in the statistics of $M_{ij}$. According to the ETH (1), in chaotic models $R_{ij}$ follow a Gaussian distribution with zero mean and variance that depends on the energies $E_{i}$ and $E_{j}$ of the two eigenstates. On the other hand, we showed that in integrable systems $M_{ij}$ is well described by a Gumbel distribution.
To investigate the effects of the non Gaussianity of $M_{ij}$, following Ref. [17], we compute the ratio $\Gamma$ defined as
@@
In Fig. 6 we plot the ratio (22) as a function of the energy difference $(E_{i}-E_{j})/\sqrt{L}$. In the main plot, we consider the matrix elements $\mathcal{O}_{ij}$ of the one-spin operator $E_{i}^{(11)}$, whereas in the inset those of the two-spin one $E^{(1,1)}_{i}E_{i+1}^{(2,2)}$ (cf. (16)). Clearly, $\Gamma$ differs significantly from $\pi/2$, for both operators. Actually, we observe $\Gamma\sim 10^{5}$ in the region with $E_{i}\approx E_{j}$. We should mention that while the finite-size scaling of $\Gamma$ in the Heisenberg chain was already investigated in Ref. [17] by employing exact diagonalization, here we provide numerical data for much larger size $L\leq 240$. The fact that $\Gamma\sim 10^{5}$ suggests that $\Gamma$ diverges in the thermodynamic limit $L\to\infty$. Indeed, we observe that $\Gamma$ grows upon increasing the total number of off-diagonal matrix elements employed to compute (22). This is compatible with the fact that $\widetilde{M}_{ij}$ is well described by a Gumbel distribution, as we now show. First, we can neglect the logarithmic correction (20) because it depends weakly on $|\mathcal{O}_{ij}|$ and it cancels out in the ratio (22). A straightforward calculation allows us to obtain from (19) the distribution of $|O_{ij}|$ as
@@
Figure 6: Value of the ratio $\Gamma$ in Eq. 22 for the operator $E_{i}^{(11)}$ for $M=20,30,40,$ (cf. (16)) and $60$ and $L=4M$ (main plot). In the inset we show the ratio $\Gamma$ for the two-spin operator $E_{i}^{(11)}E^{(22)}_{i+1}$.

### v2

@@ III.1.1 One-spin operator
Figure 5: Distribution of $M_{ij}$ (17) of the one-spin operator $E_{i}^{(11)}$ in Eq.˜16 for eigenstates of the $XXX$ chain in the finite temperature macrostate at $\beta=0.5$ with particle density $M/L=1/4$ and particle numbers $M=14,20,30$ and $40$. (a) Histogram of $\left|M_{ij}\right|$. The continuous lines are fits to the Gumbel distribution (19). In the inset we report the histograms of $\left|M_{ij}\right|/L$. (b) Histograms of the shifted $\left|M_{ij}\right|/L-\overline{\left|M_{ij}\right|}/L$, where $\overline{\left|M_{ij}\right|}$ is the average of $\left|M_{ij}\right|$. In the inset we report the averages $\overline{\left|M_{ij}\right|}$ and we compare them with the mean $\mu_{G}$ of the Gumbel distribution (19) obtained from the fits in panel (a). (c) Double logarithm of the CDF of $\left|M_{ij}\right|/L-\overline{\left|M_{ij}\right|}/L$.
In Fig.˜3 we report $M_{ij}$ (cf. (17)) for the operator $E^{(11)}_{i}$ (cf. (16)). In Fig.˜3 (a) we show the histograms $P(|M_{ij}|)$ for several chain sizes up to $L=240$, at fixed particle density $M/L=1/4$. The data are the same as in Fig.˜2 (b) for the infinite-temperature macrostate. We observe that as $L$ increases, the center of the peak moves towards larger values of $|M_{ij}|$, and the distribution broadens. Moreover, the histograms are not symmetric around the maximum. Notice that at $L=240$ the typical off-diagonal matrix elements is $M_{ij}\sim 10^{-200}$. Indeed, to correctly capture such tiny matrix elements we employed arbitrary-precision arithmetic in evaluating the ABA formulas (see Appendix A). Let us now discuss the statistics of the off-diagonal matrix elements, i.e., the functional form of the distribution $P(M_{ij})$. For the Lieb-Liniger gas, Ref. [26] showed that $P(|{M}_{ij}|/L)$ is well described by a Fréchet distribution $P_{\alpha,\beta,\nu}$, whose probability distribution function takes the form
@@
$P_{\mu,\beta}(x)=\frac{1}{\beta}\,e^{-\frac{x-\mu}{\beta}-e^{-\frac{x-\mu}{\beta}}}.$ (19)
The continuous lines in Fig. 3 are fits to (19) with $\beta,\mu$ the fitting parameters. For all the values of $L$ the data are quite well described by (19). To proceed, we observe that according to ETH (cf. (1)), one should expect that $M_{ij}\sim-L$, which has been verified in Ref. [26] for the Lieb-Liniger model. Thus, in the inset of Fig.˜3 we report the histograms $P(M_{ij}/L)$. The collapse of the data for different $L$ is not perfect, as a residual “drift” towards the right is visible. This is expected, and it was observed also for the Lieb-Liniger gas [26]. This effect can be fully accounted for by a residual $L$ dependence of the parameter $\mu$ of the distribution, which is related to the average of the Gumbel distribution $\mu_{G}=\mu+\beta\gamma$, with $\gamma$ the Euler-Mascheroni constant. Thus, let us consider the average $\overline{M_{ij}}$ as a function of $L$. Following Ref. [26] we fit $\overline{M_{ij}}/L$ to the behavior
@@
with $c,c_{0}$ fitting parameters. By fitting the data with $L\geq 80$ we obtain $c\approx 0.13$ and $c_{0}\approx 0.82$. Now, the histograms of $|M_{ij}|-c\ln(L)$ for different values of $L$ should collapse on the same curve, at least in the limit $L\to\infty$. In Fig.˜3 (b) we show the distribution of $|M_{ij}|/L-c\ln(L)$, with $c$ the fitted value. The data collapse for different $L$ improves as compared with Fig. 3 (a), although it is not perfect. This could be attributed to residual finite $L$ corrections. In the inset of Fig.˜3 (b) we report $|\overline{M_{ij}}|/L$ versus $L$ (notice the logarithmic scale on the $x$-axis). The dashed line is the fit to (20). The crosses denote the average of the Gumbel distribution $\mu_{G}$ computed from the fitted constants $\mu,\beta$ extracted in panel (a), and are compatible with $\overline{M}_{ij}/L$. It is also interesting to consider the Cumulative Distribution Function (CDF). The CDF $F(x)$ of the Gumbel distribution is
$F(x)=\int_{-\infty}^{x}P_{\mu,\beta}(t)dt=e^{-e^{-\frac{x-\mu}{\beta}}}.$ (21)
@@
Eq. (21) implies that the double logarithm of $F(x)$ is a straight line. In Fig.˜3 (c) we show the double logarithm of the CDF of $\left|M_{ij}\right|/L-c\ln(L)$ as extracted from the numerical data. The data exhibits an approximate linear behavior as a function $|M_{ij}|/L-c\ln(L)$ in the central region, whereas deviations from linear scaling are visible at the edges, which correspond to the tails of the distribution, where finite $L$ effects are expected to be larger.
@@
Figure 6: Distribution of $M_{ij}$ in Eq.˜17 of the two-spin operator $E_{i}^{(11)}E_{i+1}^{(22)}$ for $L=20,24,32,40,$ and $48$ and $M=L/4$. (a) Histogram of the absolute value $|M_{ij}|$. The continuous lines are fits to the Gumbel distribution (19), with $\mu,\beta$ fitting parameters. (b) Histograms of $\left|M_{ij}\right|/L-c\ln(L)$, with $c$ obtained by fitting the average $\overline{M_{ij}}/L$ to the expected scaling (20). The fit is reported in the inset with the gray-dashed line. We report with the cross symbols the average of the Gumbel distributions $\mu_{G}=\mu+\beta\gamma$, with $\mu,\beta$ obtained from the fits in panel (a), and $\gamma$ the Euler-Mascheroni constant. (c) Double logarithm of the CDF of $\left|M_{ij}\right|/L-c\ln(L)$.
@@
An important question is how universal is the statistics of off-diagonal matrix elements. To investigate this aspect, we consider the infinite-temperature thermal ensemble at fixed particle density $M/L=1/2$. Numerical data are reported in Fig.˜4. The qualitative behavior is similar to that at $M/L=1/4$ (see Fig.˜3). The continuous lines in Fig.˜4 (a) are fits to the Gumbel distribution (19), and are in good agreement with the numerical data. In the inset of Fig.˜4 (a) we report the rescaled histograms of $\left|M_{ij}\right|/L$. The collapse is not perfect, similar to the case with $M/L=1/4$. We employ the same strategy as for $M/L=1/4$, i.e., fitting the average $\overline{M_{ij}}/L$ to (20), and plotting in Fig. 4 the shifted $M_{ij}/L-c\ln(L)$. The fit to the expected scaling (20) gives $c\approx 0.16,c_{0}\approx 0.94$ (see the inset in panel (b)). It is clear that the probability distribution $P({M}_{ij}-c\ln(L))$ is different than that for $M/L=1/4$. The data collapse for different $L$ is better than for $M/L=1/4$, and deviations near the peak become smaller upon increasing $L$. A fit of the histogram for $L=80$ to the Gumbel (19) gives $\beta\approx 0.29,\mu\approx 0.77$. Finally, in Fig.˜4 (c) we report the double logarithm of the numerical CDF for $\left|M_{ij}\right|/L-c\ln L$. As for $M=1/4$, in the central region the CDF exhibits a clear linear behavior, as expected.
Finally, we consider the macrostate at inverse temperature $\beta=0.5$, and particle density $M/L=1/4$ in Fig.˜5. Again we find that the qualitative behavior is similar to the infinite temperature case in Fig.˜3. In Fig.˜5 (a) we fit the distribution of $|M_{ij}|/L$ to the Gumbel distribution (19), observing a good agreement. In the inset of Fig.˜5 (a) we report the histogram of the rescaled matrix elements $\left|M_{ij}\right|/L$. As it was the case for the infinite temperature macrostate, rescaling with $L$ is not sufficient to obtain a perfect collapse. Moreover, as we show in the inset of Fig.˜5 (b), the expected logarithmic behavior of the average $\overline{|M_{ij}|/L}$ cannot be easily extracted in the finite temperature case. For this reason, in the main plot of Fig.˜5 we shift the histograms by the averages $\overline{\left|M_{ij}\right|}/L$, rather than by the fit as in Figs.˜3 and 4. We observe that after subtracting the average, the histograms present a perfect collapse for all values of $L$ except for $L=160$. The discrepancy for $L=160$ is likely due to finite-size corrections. In Fig.˜5 (c) we report the double logarithm of the numerical CDF of the subtracted $M_{ij}/L-\overline{\left|M_{ij}\right|}/L$. As for the infinite temperature macrostates, the CDF exhibits a linear behavior, as expected from the Gumbel (21).
@@ III.1.2 Two-spin operator
Let us now discuss the statistics of off-diagonal matrix elements of local operators that have support on two nearest-neighbor sites. In the following we consider the infinite-temperature macrostate at $M/L=1/4$. Moreover, since we employ the approach of Ref. [25], as stressed already we are restricted to the macrostate constructed from eigenstates of the $XXX$ chain that do not contain bound states. Moreover, since the computational cost for evaluating the ABA formulas (see appendix˜A) increases exponentially with the size of the operator support, we provide results for $L\leq 60$. We anticipate that, despite this limitation, the qualitative scenario outlined for the one-spin operators remains the same. Let us focus on the operator $E_{i}^{(11)}E_{i+1}^{(22)}=(\mathds{1}_{i}-\sigma^{z}_{i})(\mathds{1}_{i+1}+\sigma^{z}_{i+1})/4$, which is sensitive to anti-aligned spin configurations on nearest-neighbor sites.
In Fig.˜6 (a) we report the histogram of $M_{ij}$. The continuous lines are fits to the Gumbel distributions (19), which are in perfect agreement with the numerical data. In Fig. 6 (b) we show the histograms of the shifted $M_{ij}-c\ln(L)$, with $c$ obtained by fitting the average $\overline{M_{ij}}/L$ to (20) (see the inset in panel (b)). The fit gives $c\approx 0.25,c_{0}\approx 0.15$. Despite the smaller chain sizes as compared with Figs.˜3 and 4, the data exhibit a reasonable collapse. A fit to the Gumbel distribution (19) gives $\beta\approx 0.28,\mu\approx 0$. Finally, in Fig.˜6 (c) we plot the double logarithm of the CDF of $\widetilde{M}_{ij}$. Again, the approximate linear behavior in the central region confirms that $M_{ij}$ are well described by the Gumbel distribution.
@@ III.1.3 Non-Gaussianity of off-diagonal matrix elements
Our results for off-diagonal matrix elements confirm that, despite integrability, the finite-size scaling is similar to the ETH scaling (1), even quantitatively. Precisely, the exponential decay as $e^{-L}$ is the same, except for the correction (20). The difference between integrable and non-integrable models emerges strikingly in the statistics of $M_{ij}$. According to the ETH (1), in chaotic models $R_{ij}$ follow a Gaussian distribution with zero mean and variance that depends on the energies $E_{i}$ and $E_{j}$ of the two eigenstates. On the other hand, we showed that in integrable systems the distribution of $\widetilde{M}_{ij}$ is well described by a Gumbel distribution with parameters that are indipendent of $L$. This is not compatible with the assumption of a Gaussian distribution for the matrix elements, as we now discuss.
To investigate the effects of the non Gaussianity of the matrix elements, following Ref. [17], we compute the ratio $\Gamma$ defined as
@@
In Fig.˜7 we plot the ratio (22) as a function of the energy difference $(E_{i}-E_{j})/\sqrt{L}$. In the main plot, we consider the matrix elements $\mathcal{O}_{ij}$ of the one-spin operator $E_{i}^{(11)}$, whereas in the inset those of the two-spin one $E^{(1,1)}_{i}E_{i+1}^{(2,2)}$ (cf. (16)). Clearly, $\Gamma$ differs significantly from $\pi/2$, for both operators. Actually, we observe $\Gamma\sim 10^{5}$ in the region with $E_{i}\approx E_{j}$. We should mention that while the finite-size scaling of $\Gamma$ in the Heisenberg chain was already investigated in Ref. [17] by employing exact diagonalization, here we provide numerical data for much larger size $L\leq 240$. The fact that $\Gamma\sim 10^{5}$ suggests that $\Gamma$ diverges in the thermodynamic limit $L\to\infty$. Indeed, we observe that $\Gamma$ grows upon increasing the total number of off-diagonal matrix elements employed to compute (22). This is compatible with the fact that $\widetilde{M}_{ij}$ is well described by a Gumbel distribution, as we now show. First, we can neglect the logarithmic correction (20) because it depends weakly on $|\mathcal{O}_{ij}|$ and it cancels out in the ratio (22). A straightforward calculation allows us to obtain from (19) the distribution of $|O_{ij}|$ as
@@
Figure 7: Value of the ratio $\Gamma$ in Eq.˜22 for the operator $E_{i}^{(11)}$ for $M=20,30,40,$ (cf. (16)) and $60$ and $L=4M$ (main plot). In the inset we show the ratio $\Gamma$ for the two-spin operator $E_{i}^{(11)}E^{(22)}_{i+1}$.
Conversely, it is interesting to derive the distribution of $M_{ij}/L$ under the assumption that the off-diagonal matrix elements follow a Gaussian distribution. For instance, let us first consider the case in which they have real and imaginary parts that are uncorrelated and Gaussian distributed with zero mean and variance $\sigma^{2}$. Now one obtains that the distribution of $M_{ij}/L$ is a Gumbel $P_{\mu,\beta}$ with parameters
$\mu=-\frac{\ln(\sigma^{2})}{L},\quad\beta=L^{-1}.$ (24)
Since the variance vanishes exponentially with increasing $L$ (cf. (1)), $\mu$ becomes finite in the large $L$ limit. However, one has $\beta=1/L$, and, crucially, $\beta$ does not depend on $\sigma$. This implies that in the large $L$ limit the distribution (24) becomes degenerate. This is in contrast with the results of Section III.1.1 and Section III.1.2, which provided numerical evidence that the parameter $\beta$ is finite and nonzero in the large $L$ limit. It is intriguing to observe that this result is in agreement with the general scenario that the difference between integrable and nonintegrable models is in the size of the fluctuations of matrix elements, those being larger in integrable systems. Finally, if, instead, the matrix elements obey (1) with $R_{ij}$ real and Gaussian distributed, then it is straightforward to check that the distribution of $M_{ij}/L$ is not of the Gumbel form.